### v1

# ASTERIX: Module for modelling the water flow on vegetated hillslopes

Abstract
The paper presents ASTERIX, an open source software for numerical integration of an extended Saint-Venant system of equations used as a mathematical tool to model the water flow from laboratory up to large-scale spatial domains applying physically-based principles of fluid mechanics. Many in-situ observations have shown that the plant cover plays a key roll in controlling the hydrological flux at a catchment scale. The plant roots facilitate the infiltration processes, the canopy intercept some proportion of rain, and plant stems slow down the flow. In case of heavy rains, the infiltration and interception processes cease in a short time, the remaining rainfall gives rise to the Hortonian overland flow and the flash flood is thus initiated. In this context, the following problem is also addressed in the article: how do the gradient of soil surface and the density of the plant cover influence the water dynamics in the Hortonian flow? The mathematical model and ASTERIX were kept as simple as possible in order to be accessible to a wide range of stakeholders interested in understanding the complex processes behind the water flow on hillslopes covered by plants.
The software is written in C programming language and it is free under GNU license. It was tested on a series of benchmark problems, laboratory experiments, and theoretical problems; and the results have shown a good agreement with the theoretical or measured data.
Keywords: Saint-Venant model  $\cdot$ porosity  $\cdot$ hexagonal raster  $\cdot$ numerical scheme  $\cdot$ hydrographic basin  $\cdot$ C programming  $\cdot$ open source
MSC2020: 35-04, 76-04 (Primary); 76-10, 35Q35, 74F10, 65M08 (Secondary)
ACM: G.1.8; G.4
Software availability
Name of software: ASTERIX - Water Flow Module
Developers: Stelian Ion, Dorin Marinescu, Stefan G. Cruceanu.
Email: stefan.cruceanu@ismma.ro
Year first available: 2024.
Program Language: C.
Software Required: 32 or 64 bit Linux operating system, gcc compiler, SDL libraries.
Program size: 5 MB.
Availability: free downloadable program sources at
http://www.ima.ro/software/ASTERIX_flow_2D.tar.xz
1 Introduction
Water is an essential element for life on earth, for plants, animals, human beings. Our daily comfort depends on water resources and our way of life is mainly shaped by them, too. The list of disasters over the past $50$ years is dominated by the water-related ones which account for $70\%$ of all deaths caused by natural catastrophes [1]. Climate change is affecting the hydrological cycle and increasing the frequency and intensity of the storms. Over $90\%$ of disasters are weather-related, including drought and aridification, wildfire, pollution and floods [2]. Fig. 1 illustrates a report regarding the water-related and non-water-related disasters in OECD countries [3].
Figure 1: Number of water-related and non-water-related disasters in OECD countries. Source: EM-DAT; The OFDA/CRED International Disaster Database, - Université catholique de Louvain (UCL) - CRED, Debarati Guha-Sapir - www.emdat.be, Brussels, Belgium.
Given all that, it is mandatory for all of us to understand the hydrological cycles to proper manage the water resources in order to protect our life and material well-being.
Mathematical modelling the hydrological processes is an effective way to estimate the risks regarding these water-related natural hazards. By modelling, one can make scenarios, can evaluate the level of the risk exposure, and can provide important information regarding an economical water use (e.g. to agronomists).
Modelling the hydrological phenomenon is a very challenging and difficult task with high complexity coming from the various processes involved in: surface runoff, soil erosion, precipitation, infiltration, evaporation, plant transpiration, root uptake, etc., see Fig. 2.
Figure 2: Watershed process
We can add the multitude of the factors (only a few being quantifiable) and the environmental heterogeneity (diversity of vegetation and soil, variations in altitude, curvature of the soil surface, etc.) and it becomes obvious that we cannot take everything into account in one single mathematical model. Higher complexity does not guarantee an overall higher performance; the performance can depend on the target purpose or hydrological variables, [4]. Besides this, a complex model often leads to problems of over-parameterization and equifinality [5, 6]. By all means, we do not argue against complex models, we only stress the fact that the modeller must be careful about the processes to be included in the models. The best models are the ones which give results closest to reality and at the same time require least number of parameters and model complexity, [7]. We add that a good mathematical model must be solvable and physically relevant.
The existent models cover a large spectrum of applications: rainfall-runoff, flood hazard management, dam break problems, agricultural hydrology, ecohydrology, and so on. There are many classification criteria of the models and a model can cast into one or more classes, depending on its task, scale or structure. For example, following [7], the models can be classified as metric, conceptual and physics-based ones. Another classification is considered in [8] as empirical, conceptual or physical. The continuous representation of hydrological processes is the main key to a physical model whose equations are based on real hydrological responses. The main advantages of such a model are:
$\bullet$ the spatial and temporal representations of the variables can be accomplished at very fine scales;
$\bullet$ it can be applied to ungauged catchments, provided that one can determine the parameters a priori;
$\bullet$ the effects of a catchment change can be explicitly represented, [7, 8].
In addition, there are many software products that implement mathematical models, some of them are commercial and some others are free. From the last category, we can mention [9, 10] for hydrology-related R and Python tools and resources. Among the most known models in the class of physical models, we mention SHE [11], MIKE-SHE [12], KINEROS [13], VIC [14], PRMS [15], SWASHES [16].
ASTERIX addresses the rainfall-runoff hydrological process using the Saint-Venant equations as a mathematical model. This software is mainly designed to study the combined effects induced by the presence of plant on soil surface and the variation of the gradient of the soil surface on the water dynamics. The plant cover and soil surface are modelled with two space distributed functions: the porosity and the soil altitude. The water-plant stem and water-soil frictional forces are quantified by two functions, both proportional with the square of the water velocity. The coefficients of proportionality are of parameter function type specific to plant cover and soil roughness, respectively. The soil altitude sets the geometrical configuration where the studied processes take place, while the porosity function and the frictional coefficients set the internal structure of this geometrical configuration. The space geometry and the internal structure give rise to the environmental configuration.
The novelty of adding a pointwise distributed porosity function to model the presence of plant cover allows the study of water dynamics for various scenarios with different vegetation distributions. Such a function can be also used to model the building distribution in urban flood studies [17, 18].
ASTERIX can be used from a laboratory level up to a hydrographic basin scale. One can use it for practical as well as for theoretical applications, such as:
$\succ$ problems with water distribution: determination of dry or flooded regions,
$\succ$ studies on the influence of plants on surface runoffs, estimating their role in the continuum Soil-Plant-Atmosphere,
$\succ$ propagation of floods produced by torrential rains,
$\succ$ the effects of severe floods due to deforestation,
$\succ$ studies on Dam Break Problems,
$\succ$ studies on Riemann Problems, etc.
For practical applications in environmental problems, the numerical approach chosen to solve such porous shallow-water equations must come with a good balance between the computational effort and the accuracy of the obtained solution on one hand, and between the precision of the measured data and the numerical accuracy on the other hand. Although high-order schemes provide better accuracy, they require higher computing effort. Furthermore, using such methods at the scale of hydrographic basins becomes very difficult due to the excessive volume of processed data. These are the reasons we developed and studied in [19] a simple discrete model based on a first-order numerical scheme with low algebraic calculations and reduced memory requirements. This scheme represents the main module of our software ASTERIX for modelling the water flow on vegetated surfaces.
The numerical scheme of the model behind this software was subjected to several internal and external validation tests. Tests such as steady flow over a bump, the oscillating solution of Thacker problem, and the solution to the Riemann problem of the shallow-water equations were used for the internal validation which consists in comparing the numerical results with the exact solutions of the model. For the external validation, we performed tests to compare the numerical results with laboratory measured data [20, 21]: dam break flow over a bump and in a L-shaped channel, downslope flow through rigid vegetation. A good agreement between the numerical and the exact/measured data was obtained in all these cases. A qualitative analysis emphasizing the similarity between the flow evolution on vegetated soil given by our numerical model and the behavior of the observed phenomenon was also performed to strengthen to confidence in the model.
We mention that the numerical modelling of the water flow in ASTERIX is accomplished on hexagonal rasters due to the cell adjacency and the local symmetry of the hexagonal structure. The advantages of using this type of network are detailed in [22].
This paper is organized as follows: in Section 2 we present the PDE (Partial Differential Equation) system of the Saint-Venant model with vegetation and the numerical scheme built to approximate this model and implemented in ASTERIX. The software is described in Section 3 and some numerical applications add new examples for additional validation on the method and the software used in Section 4. The last section presents final remarks and conclusions.
2 Mathematical model and numerical scheme
The water flow module of ASTERIX implements a numerical scheme to approximate the Saint-Venant equations with porosity. In order to better understand this software, we briefly present the key concepts concerning the mathematical model and numerical solutions in what follows. The reader is referred to [19] for more details.
2.1 Mathematical model
The Saint-Venant model with porosity we consider in this paper is based on mass and momentum balance equations and takes into account the variation of the soil surface, the plant cover, and the rain as water source. It reads as:
$\begin{split}\partial_{t}(\theta h)+\partial_{1}(\theta hv_{1})+\partial_{2}(\theta hv_{2})&=\mathfrak{M},\\ \partial_{t}(\theta hv_{a})+\partial_{1}(\theta hv_{a}v_{1})+\partial_{2}(\theta hv_{a}v_{2})+\theta h\partial_{a}{w}&=\mathfrak{t}^{p}_{a}+\mathfrak{t}^{s}_{a},\quad a=1,2.\end{split}$ (1)
The unknown variables of this PDE system are the water depth $h(t,\boldsymbol{x})$ and the two components $v_{1}(t,\boldsymbol{x})$ and $v_{2}(t,\boldsymbol{x})$ of the velocity $\boldsymbol{v}=(v_{1},v_{2})$. The external data and forcing functions of this model are:
(a) The altitude of the soil surface
$x^{3}=z(\boldsymbol{x}),\quad\boldsymbol{x}=(x^{1},x^{2})\in\Omega,$ (2)
where $\Omega\subset\mathbb{R}^{2}$ represents the horizontal projection of the soil surface and is assumed to be a connected bounded open set, $x^{1}$ and $x^{2}$ are the coordinates along two orthogonal horizontal axes, and $x^{3}$ is the coordinate along the vertical axis.
(b) The porosity function
[…]
This function takes into account the presence of the cover plant on the soil surface and is defined as the volume of void space between the plant stems (volume which can be filled with water) present in a unit volume. We note that $\theta=1$ for bare soil and $\theta=0$ for a complete sealant plant cover.
(c) The water source.
The free term $\mathfrak{M}$ quantifies the contribution of rain and infiltration to the water mass balance equation.
(d) The gravitational force.
Its action is quantified through
$w=g\left[z(\boldsymbol{x})+h\right],$ (4)
where $g$ represents the gravitational acceleration and $z(\boldsymbol{x})+h$ is the free water surface level.
(e) The water-soil frictional forces [23] are quantified by
$\mathfrak{t}^{s}_{a}=-\theta\alpha_{s}(h)|\boldsymbol{v}|v_{a},$ (5)
where $\alpha_{s}(h)$ is a non-negative function depending on the given soil surface. Experimental relations of Manning, Chézy, or the Darcy-Weisbach are some of the most used formulas for this coefficient in literature. The Darcy-Weisbach expression
$\mathfrak{t}^{s}_{a}=-\theta\alpha_{s}|\boldsymbol{v}|v_{a},\quad\rm{(Darcy-Weisbach),}$ (6)
has the advantage of being non-singular if the water depth becomes zero.
(f) The resistance opposed by plants to the water flow [24, 25] is quantified by
$\mathfrak{t}^{p}_{a}=-\alpha_{p}h\left(1-\theta\right)|\boldsymbol{v}|v_{a},$ (7)
where $\alpha_{p}$ is a non-negative constant depending on the type of vegetation.
The cumulative effects of the water-soil and water-plants interactions is assumed to be additive in this model, and one can thus introduce the resistance term
$\mathfrak{t}^{p}_{a}+\mathfrak{t}^{s}_{a}:=-{\cal K}(h,\theta)|\boldsymbol{v}|v_{a},$ (8)
where
${\cal K}(h,\theta)=\alpha_{p}h\left(1-\theta\right)+\theta\alpha_{s}(h)$ (9)
is the coefficient function of the frictional force of the water-soil-plant system.
2.2 Numerical scheme
The numerical scheme introduced in [19] to approximate the solution of model (1) casts into the general class of method of lines. We firstly construct a discretization of the spatial variables and differential operators only, and leave the time variable continuous. Then, we define a time discretization scheme to integrate the ODE (Ordinary Differential Equation) system previously obtained. A first-order FVM (Finite Volume Method) is used for the spatial discretization and a fractional step method for the time discretization.
The FVM method has three main ingredients:
(a) the net of the finite volume elements or the cells decomposition of the fluid flow domain $\Omega$,
(b) the cell discretization of the variables (unknowns or external data),
(c) the approximation of the gradient of the unknown functions on the interfaces of two neighboring cells.
(a) Cell decomposition.
Let $\{\omega_{i}\}_{i=\overline{1,N}}$ be an admissible polygonal partition [26] of $\Omega$,
$\Omega=\displaystyle\bigcup\limits_{i=1}^{N}\omega_{i},$ (10)
with $\sigma_{i}$ being the area of the cell $\omega_{i}$. For any given cell $\omega_{i}$, we denote
  • ${\cal N}(i)$ to be the set of all cell-indexes $j$ for which the cell $\omega_{j}$ has a common side $(i,j)$ with $\omega_{i}$,
  • $\boldsymbol{n}=(n_{1},n_{2})^{T}$ to be the unit normal vector pointing towards the outside of the boundary $\partial\omega_{i}$ of $\omega_{i}$.
(b) Cell discretization of the variables.
Any function $\psi$ on a cell $\omega_{i}$ is approximated by a constant value:
$\psi_{i}=\left.\psi(\boldsymbol{x})\right|_{\boldsymbol{x}\in\omega_{i}}.$ (11)
(c) Interface approximations.
For any common interface $(i,j)$ between the cells $\omega_{i}$ and $\omega_{j}$, we define the following approximations:
$\theta h_{(i,j)}:=\left\{\begin{array}[]{ll}\theta_{i}h_{i},&{\rm if}\;(v_{n})_{(i,j)}>0,\\ \theta_{j}h_{j},&{\rm if}\;(v_{n})_{(i,j)}<0,\end{array}\right.$ (12)
$\widehat{\theta h}_{(i,j)}:=\left\{\begin{array}[]{ll}\theta h_{(i,j)},&{\rm if}\;(v_{n})_{(i,j)}\neq 0,\\ \theta_{i}h_{i},&{\rm if}\;(v_{n})_{(i,j)}=0\;{\rm and}\;w_{i}>w_{j},\\ \theta_{j}h_{j},&{\rm if}\;(v_{n})_{(i,j)}=0\;{\rm and}\;w_{i}\leq w_{j},\end{array}\right.$ (13)
and
$\begin{array}[]{l}(v_{a})_{(i,j)}:=\displaystyle\frac{v_{a\,i}+v_{a\,j}}{2},\quad a=1,2,\\ (v_{n})_{(i,j)}:=\boldsymbol{v}_{(i,j)}\cdot\boldsymbol{n}_{(i,j)},\vspace{1mm}\\ w_{(i,j)}:=\displaystyle\frac{w_{i}+w_{j}}{2},\end{array}$ (14)
with $\boldsymbol{n}_{(i,j)}$ denoting the unitary normal to the common side of $\omega_{i}$ and $\omega_{j}$ pointing towards $\omega_{j}$.
Given the particular (flux conservative - partial) form of (1), we now introduce the following discrete operators:
$\begin{split}{\cal J}_{a\,i}(h,\boldsymbol{v}):=&-\displaystyle\sum\limits_{j\in{\cal N}(i)}l_{(i,j)}\theta h_{(i,j)}(v_{a})_{(i,j)}(v_{n})_{(i,j)},\\ {\cal S}_{a\,i}(h,w):=&-\displaystyle\frac{1}{2}\sum\limits_{j\in{\cal N}(i)}l_{(i,j)}(w_{j}-w_{i}){\widehat{\theta h}_{(i,j)}}n_{a}|_{(i,j)},\\ {\cal L}_{i}((h,\boldsymbol{v})):=&-\displaystyle\sum\limits_{j\in{\cal N}(i)}l_{(i,j)}\theta h_{(i,j)}(v_{n})_{(i,j)},\end{split}$ (15)
where $l_{(i,j)}$ is the length of the common interface $(i,j)$ between the cells $\omega_{i}$ and $\omega_{j}$. Combining the spatial discretization with the time fractional step method gives the full discrete scheme associated to (1) of the form
$\begin{split}\sigma_{i}(\theta_{i}h_{i})^{n+1}&=\sigma_{i}(\theta_{i}h_{i})^{n}+\triangle t_{n}{\cal L}_{i}((h,\boldsymbol{v})^{n})+\sigma_{i}\triangle t_{n}\mathfrak{M}_{i}(t^{n+1},h^{n+1}),\\ \sigma_{i}(\theta_{i}h_{i}v_{a\,i})^{n+1}&=\sigma_{i}(\theta_{i}h_{i}v_{a\,i})^{n}+\triangle t_{n}\left({\cal J}_{a\,i}((h,\boldsymbol{v})^{n})+{\cal S}_{a\,i}((h,w)^{n})\right)-\triangle t_{n}\sigma_{i}{\cal K}_{i}(h^{n+1})|\boldsymbol{v}_{i}^{n+1}|v^{n+1}_{a\,i},\end{split}$ (16)
for $a=1,2$, where a generic $\psi_{i}$ represents the value of $\psi$ on $\omega_{i}$ for $i\in\overline{1,N}$, $\varphi^{n}$ is the value of $\varphi$ at the moment of time $t^{n}$, $\triangle t_{n}:=t^{n+1}-t^{n}$ is the time step, ||$\cdot$|| ⋅ | is the euclidean norm, and
$\mathfrak{M}_{i}(t,h):=\mathfrak{M}(t,h_{i}),\quad\quad{\cal K}_{i}(h):={\cal K}(h_{i},\theta_{i}).$ (17)
To advance a time step, one must solve the system (16) for each cell $\omega_{i}$, which basically reduces to determining the water depth $h_{i}^{n+1}$ and velocity $\boldsymbol{v}_{i}^{n+1}$ at the time $t^{n+1}$ based on the known values $h_{i}^{n}$, $\boldsymbol{v}_{i}^{n}$, $t^{n}$ and $\triangle t_{n}$. This is accomplished in three steps:
  1. 1.
    Set $\triangle t_{n}$ subject to the restriction
    $\triangle t_{n}<\tau_{n},$
    where
    $\tau_{n}=CFL\displaystyle\frac{\phi_{\rm min}}{c^{n}_{\rm max}},$ (18)
    with
    $c_{i}=|\boldsymbol{v}|_{i}+\sqrt{gh_{i}},\quad c_{\rm max}=\max\limits_{i}\{c_{i}\},\quad\phi_{\rm min}=\min\limits_{i}\left\{\displaystyle\frac{\sigma_{i}}{\sum\limits_{j\in{\cal N}(i)}l_{(i,j)}}\right\}.$ (19)
  2. 2.
    Calculate $h_{i}^{n+1}$ using the mass balance relation of (16).
  3. 3.
    Solve the 2D nonlinear system of equations in $\boldsymbol{v}_{i}^{n}$ given by linear momentum relations of (16).
The time step limitation is imposed by the hyperbolic character of the shallow-water equations and by the positivity-preserving requirement of the water depth. $CFL$ is a number between 00 and $1$ (the Courant-Friedrichs-Lewy condition). The reader is referred to A for extended details on the last two steps.
When the software is used to study the Riemann Problem or models that develop shock type discontinuity, it is recommended to use an augmented linear moment flux ${\cal J}^{\boldsymbol{v}}_{a\,i}$ instead of ${\cal J}_{a\,i}(h,\boldsymbol{v})$:
${\cal J}^{\boldsymbol{v}}_{a\,i}:={\cal J}_{a\,i}(h,\boldsymbol{v})+\displaystyle\sum\limits_{j\in{\cal N}(i)}l_{(i,j)}\nu_{a(i,j)}(h,\boldsymbol{v}).$ (20)
This update is based on the artificial viscosity
$\nu_{a(i,j)}(h,\boldsymbol{v}):=c_{(i,j)}\mu_{(i,j)}\left((v_{a})_{j}-(v_{a})_{i}\right),$ (21)
where
$c_{(i,j)}=\max\{c_{i},c_{j}\},\quad\mu_{(i,j)}=\displaystyle\frac{2\theta_{i}h_{i}\theta_{j}h_{j}}{\theta_{i}h_{i}+\theta_{j}h_{j}}.$ (22)
3 Software description
ASTERIX is a software developed in C on a Linux operating system, being dedicated to the simulation of water flow on vegetated slopes. It is based on the previously described numerical scheme for the mathematical model (1) and uses a regular hexagonal network for the spatial discretization. SDL libraries are required for visualizing the time distribution of water layer depth or velocities. ASTERIX is distributed under GPL free software license.
The source code of ASTERIX is a collection of files, most of them containing many specific functions. These component functions can be logically grouped into three categories: Data In, Running and Data Out. One can improve the performance or add new actions by modifying the existent functions or adding new ones. We believe the source code and the accompanying User Guide contain sufficient details to allow a C advised programmer to make the desired modifications.
Data In.
The functions in this category collect the external information and set values to the internal variables and to the hyper-parameters of the program. The most important work is to process the environmental data concerning the soil altitude, plant cover density, and plant and soil characteristic parameters related to the frictional forces.
Domain geometry and the cell partition of the flow domain. The topographic information of the soil provided by a GIS (Geographic Information System) raster is ported to a regular hexagonal raster along this step. One edits the file hexa_parameters.dat in order to establish the number of hexagonal cells on the first row of the hexagonal grid. The number of lines of the hexagonal raster is then automatically calculated depending on the geometry of the GIS raster.
Note: The program also allows the option to work on a rectangular subdomain cut from the GIS raster.
The executable “convert” generates data files with information on the hexagonal raster (e.g. cell radius, cell centers, cell elevations, cell neighbors, cell interfaces, cell types [^1] In terms of belonging to the domain $\Omega$, the hexagonal raster of the soil surface contains internal cells and boundary cells, corresponding to the interior and boundary parts of $\Omega$, respectively., etc.) that are required by the main program for flow simulation. Details about the mathematical framework for porting data methods can be found in [22].
One can also artificially create the topography of a soil using some mathematical functions. In this case, the topographic information data can be directly generated on a regular hexagonal raster using the executable “generate”.
Boundary conditions. The software allows various boundary conditions (free discharge, given mass flux, impermeable border, Dirichlet, and combinations of these) and contains some examples in this sense. The boundary conditions are themselves a rather difficult challenge due to the hyperbolic type of the system of equations. Furthermore, identifying the parts of the border on which different conditions must be established represents another difficulty.
Parameters. The flow simulation program can be controlled through a series of parameters whose values can be modified in the ”parameter_control.dat”, file such as:
  1. 1.
    environmental parameters
    1. (a)
      water-soil friction coefficient;
    2. (b)
      water-plant friction coefficient;
    3. (c)
      plant density;
  2. 2.
    program hyper-parameters
    1. (a)
      total running time;
    2. (b)
      maximum time step;
    3. (c)
      the type of representation;
    4. (d)
      choosing an initial water layer or using one rain simulator;
    5. (e)
      extraction of snapshots at different moments of time;
    6. (f)
      control of color shades for representations;
    7. (g)
      the output data files we want to get at the end.
Running.
The flow is simulated via the executable “2D_flow” on the regular hexagonal grid. The program can be run with or without graphical representation. The option without graphics shortens the runtime for finding the numerical output.
Data Out.
The program is able to generate files with numerical data for certain quantities of interest at different moments of time. In particular, one can generate data files with the evolution of the water depth and velocities at certain predetermined points on the terrain.
When run with graphics, the program allows three graphical representations:
  1. 1.
    the relief only;
  2. 2.
    the evolution of the water depth distribution;
  3. 3.
    the evolution of the water velocity field.
If the option “relief only” is selected, ASTERIX allows the user to inspect the terrain in order to find some information about the hexagonal cells of the raster as the coordinates of their centers, the altitudes, etc. This information can be useful when one desires to establish monitoring points.
Our software includes the GIS files of some hydrographic basins with the help of which one can quickly test how the software works on real topographies. One can easily add new GIS data files for simulations on other topographies. The software has also several analytic functions implemented for simulating the flow for theoretical or laboratory experiments. The user with some experience in C can add new functions in the available sources.
The directory structure and the operating diagram of ASTERIX are synthesized in Table 1 and Fig. 3, respectively. Extended details can be found inside the User Guide.
Folder Content Description
GIS .asc DEM files containing the topography of some real terrains (e.g. ampoi, lipaia, paul, susita).
gis_hexa_conversion .c, .h The C sources and header files for porting the raster data of a file from GIS into a hexagonal raster.
generate_artificial_gis .$\rm{.c,\,.h,}$dat The source code and header files for generating the hexagonal raster of a soil surface defined by a mathematical function in the .dat file.
ini .dat Files with input data.
main .c, .h The main code files for the flow simulation.
out A place where the output (.dat, .png) files will be saved.
Table 1: Folder structure of ASTERIX with the key files.
Figure 3: Diagram (sketch) of the ASTERIX software
The variables and the parameters describing model (1) and their encoded names inside the software are also sketched in Table 2.
Encoded names in ASTERIX Vars &
params
Observations
Unknowns water_vector[$\cdot$] $h(t,\boldsymbol{x})$ Must be given at $t=0$.
velocities[$\cdot$][$\cdot$] $\boldsymbol{v}(t,\boldsymbol{x})$
External data elevation_vector[$\cdot$] $z(\boldsymbol{x})$ Quantities defining the terrain configuration.
theta_veg[$\cdot$] $\theta(\boldsymbol{x})$
alpha_terr $\alpha_{s}$
alpha_veg $\alpha_{p}$
Table 2: The variables and the main parameters of (1) with (8) and (9), and their encoded names in ASTERIX
4 Numerical examples
A suite of examples is presented in [19] for the purpose of validating the numerical scheme. In what follows, we add some new examples for additional testing on the method, as well as on the software used.
4.1 Flow on radial symmetric surfaces
A first test was performed on two theoretical surfaces with radial symmetry. These soil surfaces pictured on the first row of Fig. 4 are built using a parametric representation
$\left\{\begin{array}[]{l}x^{1}=r\cos{\theta}\\ x^{2}=r\sin{\theta}\\ x^{3}=f(r)\end{array}\right.$ (23)
on a annulus
$D:=\{(r,\theta)\left|\,r\in[r_{0},r_{1}],\,\theta\in[0,2\pi)\right.\},$
with $r_{0}=10$, $r_{1}=100$, and
$f(r)=\displaystyle A+s\cdot A\cos\left(\frac{\pi(r-r0)}{r_{1}-r_{0}}\right),$ (24)
with $A=10$ m, $s=-1$ for the Crater surface on the left column and $s=+1$ for the Hillock surface on the right column. The boundary conditions are of Dirichlet type for the upper part and free discharge for the bottom part of both surfaces:
$h(t,\boldsymbol{x})=h_{0},\quad\boldsymbol{v}(t,\boldsymbol{x})=-v_{0}\boldsymbol{n},\quad\forall\boldsymbol{x}\in\partial D^{\rm in},$ (25)
where $\partial D^{\rm in}$ is the upper boundary of the surface, $\boldsymbol{n}$ is the external oriented unit normal vector of $\partial D^{\rm in}$ at $\boldsymbol{x}$, and $h_{0}=0.05$ m, $v_{0}=1$ m/s. The advantage of such radial symmetric surfaces is that we can write the analytic solution of the flow in steady state [27].
The numerical solution was obtained with ASTERIX on a hexagonal raster with approximately $900000$ cells of radius $0.115$ m. A comparison between the analytic and numerical water depth and velocity distributions of the stationary solution of the water flow along a radial section is presented in Fig. 4. For the errors between the two solutions we used the formulas
$\epsilon_{h}=\frac{1}{N}\sum_{i=1}^{N}{\left|\frac{h_{n}(r_{i})-h_{a}(r_{i})}{h_{a}(r_{i})}\right|},\qquad\epsilon_{v}=\frac{1}{N}\sum_{i=1}^{N}{\left|\frac{v_{n}(r_{i})-v_{a}(r_{i})}{v_{a}(r_{i})}\right|},$ (26)
where $N$ is the number of points $r_{i}$ from $D$ at which the numerical solution is evaluated. $h_{n}$ and $h_{a}$ stand for the numerical and analytic water depth, respectively. $v_{n}$ and $v_{a}$ stand for the numerical and analytic water velocity, respectively.
The small values of the errors presented in Table 3 show there is a good agreement between the analytic and numerical solutions.
Crater type soil Hillock type soil
$\epsilon_{h}$ $0.0055$ $0.0111$
$\epsilon_{v}$ $0.0040$ $0.0018$
Table 3: Relative mean error between the numerical and analytic steady state solutions.
Crater type Hillock type
Soil surface
Refer to caption Refer to caption
Radial section Radial section
Soil surface
Refer to caption Refer to caption
Water depth [m]
Refer to caption Refer to caption
Water velocity [m/s]
Refer to caption Refer to caption
       $x$ [m]        $x$ [m]
Figure 4: The stationary solution of the water flow on two radial symmetric surfaces: crater type (first column) and hillock type (second column). Boundary conditions: $h=0.05$ m and $v=1$ m/s at the top of the surface and free drainage at the bottom.
4.2 Thacker Problem
The next 2D experiment, Thacker Problem, is a powerful test to check the ability of our numerical scheme to catch and highlight the propagation of a wet/dry front. The singularity appearing at the wet/dry contact between water and soil surfaces raises difficulties for many numerical methods, but a huge advantage of this problem is that it can be exactly solved [16, 28, 29, 30] and thus, one can make use of its analytic solution to theoretically test the developed numerical method.
Thacker Problem is basically formulated as the PDE system (1) modelling the flow dynamics over a paraboloid type surface
$z(x,y)=a(x-x_{0})^{2}+b(y-y_{0})^{2},\quad a,b>0,$ (27)
in the absence of vegetation ($\theta=1$) and mass source ($\mathfrak{M}=0$), for a soil resistance term of the form
$\mathfrak{t}^{s}_{a}=-\tau hv_{a},\quad a=1,2,$ (28)
where $\tau$ is a proportionality coefficient, and for a fluid velocity which does not depend on the space variable.
Depending on the values of some parameters, the exact solutions of this problem can be classified into oscillating, non-oscillating and mixed ones. In [19], we have considered the case of Thacker Problem with an oscillating solution and compared it with the numerical one described there and implemented in ASTERIX. In this subsection, we present a comparison between a non-oscillating solution for Thacker Problem and the numerical solution provided by ASTERIX [^2] The soil resistance term $\mathfrak{t}^{s}_{a}$ is theoretically chosen to be proportional to the water depth and velocity in Thacker Problem in order to obtain an analytic solution. Note that this form (28) is different from (6), and ASTERIX was modified accordingly.. The non-oscillating solution was obtained using $a=1.25\cdot 10^{-3}$, $b=5\cdot 10^{-3}$, $x_{0}=y_{0}=500\ {\rm m}$ in (27), and a proportionality coefficient of the friction force $\tau=0.7\ {\rm s}^{-1}$ in (28). The reader is referred to B for extended mathematical details about the analytic solution to this problem.
We consider the following initial data:
$u_{0}=v_{0}=0,\quad u^{\prime}_{0}=0.02g,\quad v^{\prime}_{0}=-0.1g,\quad w_{0}=15\ {\rm m},$ (29)
where $g=9.81\ {\rm m}\cdot{\rm s}^{-2}$ is the gravitational acceleration. Fig. 5 shows that our numerical solution (obtained on a regular hexagonal mesh with $487123$ cells of radius $0.888231\,{\rm m}$) is in accordance with the oscillating planar free water surface given by the analytic solution to the Thacker Problem.
$t=10\,{\rm s}$, $err=0.0107$ $t=30\,{\rm s}$, $err=0.0102$
Refer to caption Refer to caption
$t=70\,{\rm s}$, $err=0.0122$ $t=330\,{\rm s}$, $err=0.0045$
Refer to caption Refer to caption
Figure 5: 2D Thacker Problem: dynamics of the numerical free water surface. The pictures were obtained using gnuplot on data calculated with ASTERIX. The error between the numerical and analytic solutions is $err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$.
The evolution of the error $err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$ between the exact and numerical solution is pictured in Fig. 6.
Error
Refer to caption Refer to caption
time [s] time [s]
Figure 6: 2D Thacker Problem: time evolution of the error $err=||((h+z)^{app}-(h+z)^{ex})/(h+z)^{ex}||_{\infty}$ of discharge between the analytic and numerical solutions for two different hexagonal cell sizes: $1.65\,{\rm m}$ and $0.888\,{\rm m}$ for the left and right picture, respectively.
4.3 Simulation on Şuşiţa River basin
Şuşiţa’s hydrographic basin is the catchment area of Şuşiţa River in Vrancea County, Romania.
Unfortunately, we do not have data for the water distribution, plant cover density and measured velocity field in a hydrographic basin to compare our numerical results with. However, to be closer to reality, we have used the GIS data for an approximately $370\,{\rm km}^{2}$ soil surface included in a $64\times 28\,{\rm km}^{2}$ rectangular area. More precisely, this rectangular area, Fig. LABEL:fig_susita_foto, lies north-west of Tişiţa whose geographical coordinates are $45^{\circ}$ $50^{\prime}$ $36^{\prime\prime}$ North, $27^{\circ}$ $13^{\prime}$ $32^{\prime\prime}$ East.
We accomplished a theoretical experiment: starting with a uniform rainfall generated by a triangular model (see C for extended details) and characterized by
  • the total time $T_{d}=1000\,{\rm s}$ of the rain,
  • the maximum intensity $i_{a}=0.0732\,{\rm mm/s}$ of the rain, and
  • the time $t_{a}=250\,{\rm s}$ at which this maximum intensity is recorded,
on the entire basin and using two different uniform vegetation densities ($\theta_{1}=1.0$ and $\theta_{2}=0.97$), we ran ASTERIX on a network with $84651$ hexagonal cells of radius $41\,{\rm m}$.
The water distribution at a given moment in time is comparatively presented in Fig. 8 using a blue color gradient (the darkest blue corresponds to the cells with $\theta h\geq 4h_{0}$, where $h_{0}=i_{a}*T_{d}/(2*1000)=0.0366\,{\rm m}$ is the rainfall depth). The water velocity distribution at the same moment in time is pictured in Fig. 9 using a red color gradient for the modulus of the velocity and arrows for the flow direction. The convention of coloring the relief where $\theta h\leq 0.1h_{0}$ is applied for both figures. The data in Table 4 confirms that more water is retained in the basin if the vegetation is denser.
$\theta$ Rainfall Water out Water left
- [m3] [m3] [m3]
$1.00$ $3006.6$ $23.8$ $2982.8$
$0.97$ $3006.6$ $20.7$ $2985.9$
Table 4: Data for the theoretical experiment on Şuşiţa’s hydrographic basin. ASTERIX was run for $t=3000$ s.
Figure 8: Snapshots of water distribution on Şuşiţa’s hydrographic basin for two different uniform vegetation densities: $\theta=1.0$ and $\theta=0.97$ on the upper and bottom picture, respectively. As expected, our numerical data are consistent with terrain observations: the amount of water leaving the basin is greater in the case of lower vegetation density.
Figure 9: Snapshots of water velocity distribution on Şuşiţa’s hydrographic basin for two different uniform vegetation densities: $\theta=1.0$ and $\theta=0.97$ on the upper and bottom picture, respectively. As expected, the velocities are smaller when vegetation is denser.
Furthermore, Fig. 10 emphasizes that the maximal discharge rate is smaller and occurs later when vegetation is denser.
Rainfall rate [mm/s]
Discharge rate [m3/s]
Time [s]
Figure 10: Simulation of a rainfall and the discharge rates on Şuşiţa’s hydrographic basin for two different uniform vegetation densities ($\theta=0.97$ and $\theta=1$).
Remark 4.1.
One can easily test ASTERIX (without having to modify the code) for a terrain given by a GIS data file (e.g. susita.asc). Assuming that the source code is already compiled, launch from the ’main’ directory
\$ ./natural_relief_selection.sh susita
\$ ./2D_flow
The reader is referred to the User Guide for extended details.
4.4 Sensitivity
One of the main issues in mathematical modelling of real life phenomena is the sensitivity of the solution of the model to the variation of the model parameters. Roughly speaking, the sensitivity analysis highlights how large is the variation of the solution with respect to the variation of the parameters. There is a large variety of methods dedicated to sensitivity analysis; some of them perform very well (in the sense that they offer relevant results) for a class of models, but they become inadequate for other classes. A very rich literature is devoted to this subject, see [32, 33], for example, for a comprehensive view of the sensitivity analysis concepts and methods.
Regarding the vegetated Saint-Venant model, we consider the water-soil $\alpha_{s}$ (as in the Darcy-Weisbach law) and the water-plant $\alpha_{p}$ friction coefficients as the parameters of the model. The porosity $\theta(\boldsymbol{x})$ and the altitude $z(\boldsymbol{x})$ functions identify a physical configuration and are considered to be time independent. The dependence of the hydro-dynamical variables on the physical configuration is obvious and will not be the subject of sensitivity analysis; it is rather a problem of ability of the vegetated Saint-Venant model and numerical ASTERIX software to offer a reliable solution to a given hydrological problem. We note that our analysis is only a sketch, a systematic analysis being beyond the scope this article.
Estimation of friction coefficients
The parameters $\alpha_{s}$ and $\alpha_{p}$ can not be directly measured, they must be estimated by other methods. Therefore, it is very important to know how its imprecise assessment affects the solution of the Saint-Venant equation.
To find these parameters, we use here an inverse method and the measurements of an experiment reported in [21]. The experimental installation consists of a rectangular laboratory flume ($18$ m long and $1$ m wide) with a longitudinal bottom slope $S=1.05$ mm/m and partially covered with vegetation. The vegetation is modelled using uniformly distributed emergent circular cylinders of radius $R_{c}=5\,{\rm mm}$ and has a density of $N=81\,{\rm cylinders}/{\rm m}^{2}$. This type of vegetation allows the estimation of the porosity by $\theta_{v}=1-N\pi R_{c}^{2}\approx 0.99364$. The experiment was performed for different given values of the upstream flow rate $q_{L}$, free downstream discharge, two types of vegetation distribution,
$\theta_{1}=\left\{\begin{array}[]{ll}1,&x\in[0,9)\\ \theta_{v},&x\in[9,18]\end{array}\right.,\quad\theta_{2}=\left\{\begin{array}[]{ll}\theta_{v},&x\in[0,9)\\ 1,&x\in[9,18]\end{array}\right.$
and $z=-m(x-18),\,x\in[0,18]$, with $m=0.105$ mm/m. The numerical simulation is carried out using ASTERIX with the following boundary conditions:
$uh(t,0)=q_{L},\quad\partial_{x}u(t,18)=\partial_{x}h(t,18)=0.$
We will perform later a sensitivity local analysis around of some reference values $(\alpha^{R}_{s},\alpha_{p}^{R})$ determined by fitting the experimental data reported in [21]. To illustrate the difficulties of estimating these friction parameters, we minimize the errors between the measured and calculated water depth on a $21\times 26$ grid inside the domain
$\Lambda:=\displaystyle\{(\alpha_{s},\alpha_{p})\big{|}\quad 0.0025\leq\alpha_{s}\leq 0.0185,\quad 55\leq\alpha_{p}\leq 80\}.$
In this sense, we define the error function
$\chi(\alpha_{s},\alpha_{p}):=\sqrt{\frac{1}{N_{e}N_{g}}\sum_{l=1}^{N_{e}}\sum_{i=1}^{N_{g}}\displaystyle\left(h^{\rm exp}(x_{i}^{l})-h(x_{i}^{l};\alpha_{s},\alpha_{p})\right)^{2}},$ (30)
where $N_{e}$ is the number of considered experiments, $N_{g}$ is the number of gauges placed inside the channel for measurements, $x_{i}^{l}$ is the position of the gauge $G_{i}$ on the $l^{\rm th}$ experiment, while $h^{\rm exp}$ and $h$ are the measured and calculated water depths, respectively. For example, if all six experiments (with $N_{g}=17$ gauges) reported in [21] are used, then the error takes the following form:
$\chi(\alpha_{s},\alpha_{p})=\sqrt{\frac{1}{6\cdot 17}\sum_{j,k,i}\displaystyle\left(h^{\rm exp}(x_{i}^{k};q_{jk};\theta_{k})-h(x_{i}^{k};q_{jk};\theta_{k},\alpha_{s},\alpha_{p})\right)^{2}}.$ (31)
with
$j=\overline{1,3},\quad k=\overline{1,2},\quad i=\overline{1,17},$
and
$q_{11}=7\,{\rm l/s},\quad q_{21}=15\,{\rm l/s},\quad q_{31}=21\,{\rm l/s},$
$q_{12}=7\,{\rm l/s},\quad q_{22}=15\,{\rm l/s},\quad q_{32}=50\,{\rm l/s}.$
Fig. 11 presents the contour levels of the error surface for the following cases:
Case I: Experiment with $\theta=\theta_{1}$, $q=q_{21}$;
Case II: Experiments with $\theta=\theta_{1}$, $q\in\{q_{11},q_{21},q_{31}\}$;
Case III: Experiments with $\theta=\theta_{2}$, $q\in\{q_{12},q_{22},q_{32}\}$;
Case IV: All six experiments: $\theta=\theta_{1}$, $q\in\{q_{11},q_{21},q_{31}\}$ and $\theta=\theta_{2}$, $q\in\{q_{12},q_{22},q_{32}\}$.
Case I Case II
Plant friction: $\alpha_{p}$
Refer to caption Refer to caption
Case III Case IV
Plant friction: $\alpha_{p}$
Refer to caption Refer to caption
Soil friction: $\alpha_{s}$ Soil friction: $\alpha_{s}$
Figure 11: The contour level lines of the error surface in the four analyzed cases
When using only one experiment, e.g. Case I, the minimum is poorly located: there are many pairs $(\alpha_{s},\alpha_{p})$ in the area of minimum error and it is not possible to decide which one is the best. For Case II and Case III, each with 3 experiments, the parameters are not yet sufficiently well located: $\alpha_{s}$ is poorly located in Case II and $\alpha_{p}$ is poorly located in Case III. For Case IV, the surface of the error function (31) is drawn in Fig. 12 and the minimum is the better located.
Figure 12: The error surface for Case IV. The minimum of this function is reached at $\alpha_{s}=0.00709,\,\alpha_{p}=73.39$: $\quad{\rm min}=\chi\left(0.00709,73.39\right)=2.26$ mm.
For solving a minimization problem, more efficient methods can of course be used than searching for the solution on a grid. Since the evaluation of the cost function (31) associated with all six experiments at one single point requires a considerable computational effort, we used the “Optimization by pattern search” algorithm [34] based on the “Direct Search” method introduced in [35] and found
$\alpha_{s}=0.00709\quad{\rm and}\quad\alpha_{p}=73.39.$ (32)
A comparison between the measured data and numerical water depth found with these determined parameter values is illustrated in Fig. 13.
     $\theta=\theta_{1}$   $\theta=\theta_{2}$
Water depth [mm]
Refer to caption Refer to caption
Water depth [mm]
Refer to caption Refer to caption
Water depth [mm]
Refer to caption Refer to caption
        x [m]         x [m]
Figure 13: Longitudinal profiles of water depth for the flow over a slope. The vegetated and the bare parts of the soil are marked with green and white, respectively.
For each experiment, the error $\epsilon=\chi\left(0.00709,73.39\right)$ (in [mm]) between the two solutions is calculated with (30) for $N_{e}=1$ and $N_{g}=17$; Table 5 presents the values of these errors.
$\theta$ $\theta_{1}$ $\theta_{2}$
$q$ $q_{11}$ $q_{12}$ $q_{13}$ $q_{21}$ $q_{22}$ $q_{23}$
$\epsilon$ 1.48 2.05 1.73 3.08 2.10 2.73
Table 5: The errors between the numerical solution and the experimental data pictured in Fig. 13
Sensitivity with respect to the discharge $q$, porosity $\theta_{v}$ and slope $m$
We further proceed and make a theoretical analysis for the response of the averaged water depth value along the vegetated zone $9\leq x\leq 18$ to the variation of the environmental variables $q$, $\theta_{v}$ and $m$. Let us introduce $\overline{h}\left((\alpha_{s},\alpha_{p});(q,\theta_{v},m)\right)$ by
$\overline{h}\left((\alpha_{s},\alpha_{p});(q,\theta_{v},m)\right):=\frac{1}{N_{v}}\sum_{i}h\left(x_{i};(\alpha_{s},\alpha_{p});(q,\theta_{v},m)\right),$ (33)
where $N_{v}$ is the number of grid points $x_{i}\geq 9$ taken inside the vegetated zone, $h\left(x_{i};(\alpha_{s},\alpha_{p});(q,\theta_{v},m)\right)$ is the steady water depth value given by the numerical model evaluated at $x_{i}$ with the following environmental variables: discharge $q_{L}=q$, the porosity $\theta_{v}$ and the slope $m$. One defines a regular grid inside the local domain $\Lambda$:
$\alpha^{i}_{s}=\underline{\alpha_{s}}+(i-1)\triangle_{s},\quad\alpha^{j}_{p}=\underline{\alpha_{p}}+(j-1)\triangle_{p},\quad i=\overline{1,N_{s}},\quad j=\overline{1,N_{p}}.$
The sensitivity is quantified by:
$\begin{split}\delta_{s}\overline{h}(\boldsymbol{p})&:=\displaystyle\frac{1}{(N_{s}-1)N_{p}}\sum_{i=1}^{N_{s}-1}\sum_{j=1}^{N_{p}}\displaystyle\frac{\overline{h}((\alpha^{i+1}_{s},\alpha^{j}_{p});\boldsymbol{p})-\overline{h}((\alpha^{i}_{s},\alpha^{j}_{p});\boldsymbol{p})}{\triangle_{s}},\\ \delta_{p}\overline{h}(\boldsymbol{p})&:=\displaystyle\frac{1}{N_{s}(N_{p}-1)}\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{p}-1}\displaystyle\frac{\overline{h}((\alpha^{i}_{s},\alpha^{j+1}_{p});\boldsymbol{p})-\overline{h}((\alpha^{i}_{s},\alpha^{j}_{p});\boldsymbol{p})}{\triangle_{p}},\end{split}$ (34)
where $\boldsymbol{p}$ stands for the environmental variables $\boldsymbol{p}=(q,\theta_{v},m)$. The two values of these functions $\delta_{s}$ and $\delta_{p}$ for a given environmental variable $\boldsymbol{p}$ represent some average measures of the variation of the water depth $\overline{h}$ defined in (33) with respect to the water-soil $\alpha_{s}$ and water-plant $\alpha_{p}$ friction coefficients, respectively.
Fig. 14 illustrates the behavior of the sensitivities $\delta_{s}$ (top row) and $\delta_{p}$ (bottom row) of the model with respect to the environmental variables $\boldsymbol{p}$. The porosity $\theta$ is fixed to $\theta_{v}=0.9936$ on the first column and the slope $m$ is fixed to the value $0.105$ on the second column. A first observation we can make is that, in general, the sensitivity increases with the discharge $q$. An exception can be noticed for $\delta_{s}$ at small values of $\theta_{v}$, see the picture B. As expected, a lower vegetation density leads to higher sensitivity with respect to the water-soil friction coefficient and lower sensitivity with respect to the water-plant friction coefficient, see pictures B and D. Also, as the longitudinal slope $m$ increases (the gravitational forces start to dominate), the sensitivity with respect to the water-plant friction coefficient diminishes, see picture C. We can also observe that the sensitivity with respect to the water-soil friction coefficient increases for small values of $m$ and decreases for higher values of $m$, see picture A.
A B
$\delta_{s}$
Refer to caption Refer to caption
C D
$\delta_{p}$
Refer to caption Refer to caption
$m\,[\%]$ $\theta_{v}$
Figure 14: Sensitivities $\delta_{s}$ (top row) and $\delta_{p}$ (bottom row) of the model with respect to the environmental variables $\boldsymbol{p}=(q,\theta_{v},m)$. The values of the sensitivities on the first column are calculated with constant porosity $\theta_{v}$. The values of the sensitivities on the second column are calculated with constant slope $m$.
5 Conclusions and final remarks
This paper describes a new module of ASTERIX dedicated to compute the solution of Saint-Venant equations in the presence of vegetation for water flow modelling. Based on a first-order numerical scheme, the computational effort is reasonable and this makes it suitable to be used from laboratory level up to practical applications for environmental problems where the flow surface can be quite large.
The software is designed to work on regular spatial hexagonal network. When the terrain data are provided by a GIS file, the software uses the “porting data” module [22] which transfers the information to a hexagonal raster. It can also directly generate information (topography, vegetation, etc) on hexagonal rasters for theoretical or laboratory experiments. For modelling the water flow, one can start with an initial water layer or use a rain simulator (e.g. a triangular hyetograph as in C).
To run ASTERIX, one must know the values of some environmental parameters in order to get results as close to the real phenomena as possible. For example, the friction coefficients $\alpha_{s}$ and $\alpha_{p}$ cannot be directly measured; instead, one can use some reference values from literature or they can be evaluated through some inverse methods using laboratory measurements to fit the data, see Subsection 4.4 as a sketch example. Note that the estimation of these friction parameters is influenced by the distribution of the vegetation. For example, it is more difficult to locate $\alpha_{s}$ on the flume experiments with downstream vegetation (Fig. 11, Case II) and to locate $\alpha_{p}$ when the vegetation is placed upstream (Fig. 11, Case III). Fig. 11, Case I emphasizes that one can find a large number of pairs $(\alpha_{s},\alpha_{p})$ leading to very small errors and that it is difficult to decide the correct parameters when only one experiment is used. Their best localization is obtained when using all the experiments (Fig. 11, Case IV). We can also imagine other scenarios (with variable $\theta$ or different values of the slope $m$, for example), but we did not go further because we wanted to compare the results with experimental measurements. This investigation suggest that one must cleverly choose a set of experiments with different configurations in order to determine the parameters as best as possible.
In real conditions, the terrain is heterogeneous and working only with average values for $\alpha_{s}$ and $\alpha_{p}$ can be risky and, therefore, distributed values on the entire domain should be used instead. In such cases, a possible solution is to partition the surface into quasi-homogeneous subdomains with constant friction parameters.
It should be noted that, in addition to the distribution of the plant cover, the configuration of the soil surface and the water discharge for which the measurements are made are important for the determination of $\alpha_{s}$ and $\alpha_{p}$. This is also highlighted by the theoretical investigation of the sensitivity of the model with respect to $q$, $\theta_{v}$ and $m$ from the second part of Subsection 4.4, as can also be noted from Fig. 14.
Future Extensions
$\circ$ ASTERIX does not yet take into account the water infiltration and interception, but an extension in this sense is in progress. For example, adapting the mass sources $\mathfrak{M}$ through algebraic type laws can be a solution to include these processes.
$\circ$ A module for the erosion and sedimentation processes based on the Hairsine-Rose model [36, 37, 38] is being prepared to be attached to our software in the near future.
$\circ$ Some spatial distributed quantities are only available as scattered data affected or not by noise. In this sense, we are also considering implementing a module for treating and porting such data on hexagonal raster.
Appendix A Details on the Numerical Scheme
The numerical scheme (16) is the result of combining the following systems from [19]
$\sigma_{i}(\theta_{i}h_{i})^{*}$ $=\sigma_{i}(\theta_{i}h_{i})^{n}+\triangle t_{n}{\cal L}_{i}((h,\boldsymbol{v})^{n}),$ (35a)
$\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{*}$ $=\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{n}+\triangle t_{n}\left({\cal J}^{v}_{a\;i}((h,\boldsymbol{v})^{n})+{\cal S}_{a\;i}((h,w)^{n})\right),$ (35b)
$\sigma_{i}(\theta_{i}h_{i})^{n+1}$ $=\sigma_{i}(\theta_{i}h_{i})^{*}+\sigma_{i}\triangle t_{n}\mathfrak{M}_{i}(t^{n+1},h^{n+1}),$ (36a)
$\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{n+1}$ $=\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{*}-\triangle t_{n}{\cal K}_{i}(h^{n+1})|\boldsymbol{v}_{i}^{n+1}|v^{n+1}_{a\;i}.$ (36b)
with $a=1,2$. To get the solution of (16), one must solve
$\bullet$ the scalar nonlinear equation for $h^{n+1}_{i}$ obtained by combining (35a) and (36a), and
$\bullet$ the 2D nonlinear system of equations for velocity $\boldsymbol{v}^{n+1}_{i}$ obtained by combining (35b) and (36b).
This second step has an analytic solution of the form
$\boldsymbol{v}^{n+1}_{i}=\mu\cdot(\theta_{i}h_{i}\boldsymbol{v}_{i})^{*},$ (37)
where
$\mu=\frac{2}{\alpha+\sqrt{\alpha^{2}+4\beta|\boldsymbol{\gamma|}}},$ (38)
with
$\alpha:=(\theta_{i}h_{i})^{n+1},\quad\beta:=\triangle t_{n}{\cal K}(h_{i}^{n+1},\theta_{i}),\quad\gamma_{a}:=(\theta_{i}h_{i}v_{a\;i})^{*}.$
Note that for the particular cases when the mass source $\mathfrak{M}$ does not depend on $h$ (e.g. only the rain is taken into consideration), the computation of $h^{n+1}_{i}$ is direct and does no longer require to solve a nonlinear equation.
Appendix B Analytic Solution for Thacker Problem
We present here a particular analytic solution of the shallow-water equations (1) in the absence of vegetation ($\theta=1$), for $\mathfrak{t}^{p}=\boldsymbol{0}$ and $\mathfrak{t}^{s}=h\tau\boldsymbol{v}$, where $\tau$ is a proportionality coefficient, for a soil surface $z(\boldsymbol{x})$ of the form
$z(x,y)=a(x-x_{0})^{2}+b(y-y_{0})^{2},\quad a,b>0,$ (39)
and for a water velocity which does not depend on space variable, i.e. $\boldsymbol{v}(t,\boldsymbol{x})=\boldsymbol{f}(t)$. To avoid confusions in formulas, we consider
$\boldsymbol{x}=(x,y)^{T}\quad{\rm and}\quad\boldsymbol{v}=(u,v).$ (40)
A damped non-oscillating solution is obtained if
$\Delta_{a}:=\tau^{2}-8ga>0\quad{\rm and}\quad\Delta_{b}:=\tau^{2}-8gb>0.$ (41)
In order to write the general solution of the flow, we introduce the notations:
$u_{0}:=u(0),\quad u^{\prime}_{0}:=u^{\prime}(0),\quad v_{0}:=v(0),\quad v^{\prime}_{0}:=v^{\prime}(0),$ (42)
$\lambda_{1}=\lambda_{1}^{\Delta}:=\frac{-\tau-\sqrt{\Delta}}{2},\quad\lambda_{2}=\lambda_{2}^{\Delta}:=\frac{-\tau+\sqrt{\Delta}}{2},$ (43)
$A=A^{\Delta}:=\frac{V_{0}\lambda_{2}-V^{\prime}_{0}}{\sqrt{\Delta}},\quad B=B^{\Delta}:=-\frac{V_{0}\lambda_{1}-V^{\prime}_{0}}{\sqrt{\Delta}},$ (44)
$V(t;\Delta,V_{0},V^{\prime}_{0}):=Ae^{\lambda_{1}t}+Be^{\lambda_{2}t},$ (45)
$W_{0}(t;\Delta,V_{0},V^{\prime}_{0}):=\frac{1}{2g}\left[A^{2}\frac{\lambda_{2}}{\lambda_{1}}\left(1-e^{2\lambda_{1}t}\right)+B^{2}\frac{\lambda_{1}}{\lambda_{2}}\left(1-e^{2\lambda_{2}t}\right)+2AB\left(1-e^{-\tau t}\right)\right]$ (46)
$\begin{split}w_{0}(t):=&w_{0}(t;\Delta_{a},\Delta_{b},u_{0},v_{0},u^{\prime}_{0},v^{\prime}_{0})\\ :=&w_{0}+W_{0}(t;\Delta_{a},u_{0},u^{\prime}_{0})+W_{0}(t;\Delta_{b},v_{0},v^{\prime}_{0}),\end{split}$ (47)
The general solution can now be written as
[…]
where $[\alpha]_{+}$ is the positive part of $\alpha$ and
$w(t,x,y):=w_{0}(t)-\frac{1}{g}(x-x_{0})(\tau u(t)+u^{\prime}(t))-\frac{1}{g}(y-y_{0})(\tau v(t)+v^{\prime}(t))$ (49)
is the evolution of the free water surface.
Note: A damped oscillating solution is obtained if both inequalities (41) are not satisfied; if only one of the inequalities is not satisfied, then the solution is oscillating only in one direction.
Appendix C Rainfall Hyetograph: triangular model
This rainfall model has three parameters: the total time $T_{d}$ of the rain, the maximum intensity $i_{a}$ of the rain and the time $t_{a}$ at which this maximum intensity is recorded. The ratio between the total time $T_{d}$ and the moment of time $t_{a}$ is called the rainfall advance coefficient.
$\gamma=\frac{t_{a}}{T_{d}}.$
The graph of the instantaneous rain intensity $r(t)$ over the time interval $[0,T_{d}]$ has the shape of a triangle with $T_{d}$ and $i_{a}$ as its base and height, respectively (see Fig. 15).
Figure 15: Hyetograph of a rainfall: triangular model.
$r(t)=\left\{\begin{array}[]{ll}i_{a}\displaystyle\frac{t}{t_{a}},&0\leq t\leq t_{a},\\ i_{a}\displaystyle\frac{1-t/T_{d}}{1-\gamma},&t_{a}<t\leq T_{d},\\ 0,&{\rm otherwise}.\end{array}\right.$
The amount of water falling over the interval $[t_{1},t_{2}]$ is:
$V(t_{1},t_{2})=\int\limits_{t_{1}}^{t_{2}}r(t){\rm d}t,$
and therefore, the total amount of water falling over the interval $[0,T_{d}]$ is equal to the area of the triangle:
$V=\frac{i_{a}T_{d}}{2}.$